how can i use the ac test to determine which of the following trinomials can be factored and for each one how can i find m and n for each trinomials that can be factored.
1. 2p-6q+pq-3q^2
2. m^4-9n^4

Answer by stanbon(75887)

1. 2p-6q+pq-3q^2
Factor the first two and then the last two terms separately.
=2(p-3q)+q(p-3q)
Factor out the common factor:
=(p-3q)(2+q)
---------------
2. m^4-9n^4
=(m^2+3n^2)(m+3n)(m-3n)
-------------
Cheers,
Stan H.